# Lieb-Schultz-Mattis Anomalies and Anomaly Matching

### 01 Apr 2026 v1

@@ 4.4 Topological Quantum Spin Liquids
For instance, a $\mathbb{Z}_{2}$ TQSL has 4 types of anyons, often denoted by 1, $e$, $m$ and $\epsilon$, where 1 is a trivial anyon that is similar to the conventional magnons, $e$ and $m$ are both anyons with bosonic self-statistics but semionic mutual braiding statistics [^7] That two particles have semionic mutual braiding statistics means that when one particle moves around the other by a full circle, the many-body state will get a 1-1 phase factor., and $\epsilon$ can be viewed as the bound state of $e$ and $m$, which has fermionic self-statistics and semionic mutual braiding statistics with both $e$ and $m$ [73]. If the $\mathbb{Z}_{2}$ TQSL further has an $\mathrm{SO}(3)$ spin rotational symmetry, then nontrivial anyons $e$, $m$ and $\epsilon$ may carry half-integer-spin projective representation of the $\mathrm{SO}(3)$ symmetry, instead of only carrying integer-spin linear representations as do magnons. Anyons carrying half-integer spins is an example of symmetry fractionalization, and different patterns of symmetry fractionalization signal different symmetry-enriched $\mathbb{Z}_{2}$ TQSLs [74, 56].
@@
Below we discuss a concrete example, i.e., $\mathbb{Z}_{2}$ TQSL on triangular lattice spin-1/2 systems with $G_{\rm UV}=p6m\times\mathrm{SO}(3)\times\mathbb{Z}_{2}^{\mathsf{T}}$, first worked out in Refs. [79, 80]. In this context, we only need the fractionalization patterns of the $G_{\rm UV}$ symmetry on the $e$ and $m$ anyons, which can be largely understood using general arguments, without doing complicated calculations.
@@
We can further fix the fractionalization pattern of the $p6m$ symmetry on $m$ completely. To this end, imagine inserting an $\mathrm{SO}(3)$ monopole into the system, which can be viewed as a $2\pi$ flux of the $\mathrm{U}(1)$ spin rotation symmetry around the $z$ axis. As the $2\pi$ flux is gauge-equivalent to no flux, this process nucleates an anyon $v$. The fact that $e$ ($m$) carries half-integer (integer) spin means that $e$ ($m$) experiences a 1-1 (1) Aharonov-Bohm phase factor when circling around this flux, which uniquely fixes the braiding phase between $e$ ($m$) and $v$, and further fixes $v=m$. This argument is essentially Laughlin’s flux-threading argument in fractional quantum Hall effect.
Finally, note that the discussion in Sec. 4.2 shows that the projective representation of the $p6m$ symmetry carried the $\mathrm{SO}(3)$ monopole is completely fixed by the type of the lattice. Therefore, the projective representation under the $p6m$ symmetry carried by $m$ is also completely fixed by the type of the lattice, which can be read off from Sec. II B 1 in Ref. [28]. For example, for $m$ translations along the two directions do not commute.
Put together, the above discussion shows that the symmetry fractionalization pattern of the entire $G_{\rm UV}=p6m\times\mathrm{SO}(3)\times\mathbb{Z}_{2}^{\mathsf{T}}$ symmetry on $m$ is fully fixed, and $e$ carries half-integer spin under $\mathrm{SO}(3)$ and Kramers doublet under $\mathbb{Z}_{2}^{\mathsf{T}}$. However, the fractionalization pattern of the $p6m$ symmetry on $e$ has not been determined. To determine it, one way to proceed is to perform the systematic anomaly-matching analysis in Ref. [85]. A general framework to calculate the pullback of the IR anomaly $\varphi^{*}(\omega_{\rm IR})$ of a symmetry-enriched TQSL is given in Ref. [86], then using the UV anomaly obtained in Ref. [28], one can classify all possible symmetry-enrichment patterns compatible with the anomaly-matching condition Eq. (18). We refer the readers to these references for more details, as they are relatively technical. Alternatively, one can also derive the symmetry fractionalization pattern of $e$ using the physical picture that the background anyon ($e$ here) must exhibit the same fractionalization pattern of space group as the DOF in the unit cell (see Ref. [80]).
@@ 4.5 Compressible Phases
As our last example, we briefly review the matching of filling anomaly in compressible phases. Here we consider a system with $\mathrm{U}(1)$ and translation symmetries, and a compressible phase is one where the low-energy effective field theory of the system takes the same form at different filling fractions. The most familiar example are superfluids and Fermi liquids. Being compressible is a rather strong condition. For example, Dirac and Weyl semimetals are not compressible in the above sense, because when the filling fraction changes, they will become Fermi liquids and the effective theory changes.
By matching the filling anomaly, Ref. [87] pointed out that, in spatial dimension larger than 1, the internal symmetry of the effective theory of a compressible phase cannot be a 0-form symmetry described by a finite-dimensional compact Lie group. To get some intuition of this result, note that the effective theory of a $(d+1)$-dimensional superfluid actually has a (d1)(d-1)-form symmetry, which corresponds to the conservation of vorticity [55, 88]. In Fermi liquids, the symmetry of the effective theory is a “loop group” of the $\mathrm{U}(1)$, which corresponds to the conservation of fermion number at each point of the Fermi surface [87]. In both cases, the symmetries of effective theories are not 0-form symmetry described a finite-dimensional compact Lie group. We refer the readers to Refs. [87, 89] for the reasoning behind the above result.
@@ 5 OTHER GENERALIZATIONS
In this section, we discuss various other generalizations of the LSM anomalies. We remark that there are generalizations not covered in this review due to the limit of length, and we refer the readers to Refs. [21, 22, 23, 24, 25, 26] for details.
@@ 5.1 Disordered LSM and Average Anomaly
In real solids, lattice symmetries are always broken by impurities. However, viewed as an ensemble (assuming self averaging), lattice symmetries like translations are usually restored “on average”. That is, while the symmetry is completely broken for any given disorder realization, in the ensemble any two disorder realizations related to each other via translations occur with the same probability. Therefore, it is important to understand to what extent the LSM constraints apply in this case. Remarkably, LSM anomalies can still be robust if the lattice symmetries are only preserved on average [46, 11, 90, 91].
As a concrete example, consider a spin chain with an average translation symmetry, which still allows one to define a unit cell. Moreover, suppose there is an on-site $\mathrm{SO}(3)$ spin rotational symmetry such that each unit cell contains a spin-1/2 moment. If the translation symmetry was exact, in Sec. 2 we see that this symmetry setting would lead to a nontrivial LSM anomaly that forces all symmetric states to be LRE. Interestingly, when the translation symmetry is only preserved on average in an ensemble of such systems [^8] By the Imry-Ma argument, such an average translation symmetry cannot be broken spontaneously in 1D., then the probability for $\mathrm{SO}(3)$ symmetric ground states to be LRE approaches 1 in the thermodynamic limit [11]. That is, the ensemble is LRE. One can further establish an analog of Theorem 2.3 that bounds the decay of ensemble-averaged spin correlation function [46].
@@ 5.2 Fermionic Systems
A physically clear example is spin-1/2 electrons at half filling (i.e., one electron per unit cell). Suppose there is a large Hubbard-like interaction suppressing double occupancy, then the effective description of the system is a spin-1/2 system satisfying the previously discussed LSM theorem. It is then natural to expect that similar constraint applies even when the single-occupancy constraint is not strictly enforced. That is, the system cannot be in a symmetric SRE state at half filling.
@@ 5.3 SPT-LSM Theorems
Other examples of SPT-LSM theorems are considered in Refs. [17, 18, 19]. For the general theory for SPT-LSM theorems, we refer the readers to Refs. [20, 8].
@@ 6 CRYSTALLINE SYMMETRY-PROTECTED TOPOLOGICAL PHASES
Finally, we discuss a fruitful approach to characterize ’t Hooft anomalies, the “anomaly in-flow”. In this approach, the $d$-dimensional system with the anomalous symmetry is viewed as the boundary of a $(d+1)$-dimensional bulk, such that the whole system has no anomaly. That is, the non-gauge-invariant response of the boundary system to the background gauge field is cancelled by the response of the bulk. To achieve this, the bulk has to be a nontrivial SPT phase. This picture can be made precise in the QFT framework, establishing a one-to-one correspondence between ’t Hooft anomalies in $d$ dimensions and SPT phases in $(d+1)$ dimensions.
This anomaly in-flow picture can also be applied to LSM anomalies, where the bulk is an SPT phase protected by internal and/or crystalline symmetries (hence “crystalline SPTs”) [27, 94]. As a concrete example, let us revisit the LSM anomaly for the spin-1/2 chain discussed in Sec. 2.1. The 2D bulk is easily seen to be a decoupled stack of Haldane chains (i.e., the 1D nontrivial SPT phase protected by $\mathrm{SO}(3)$ symmetry), each, say, along the $y$ directions. Moreover, these Haldane chains have a translation symmetry along the $x$ direction perpendicular to $y$. The spin-1/2 chain emerges at the edge along $x$ direction. Similar constructions apply to other LSM anomalies discussed in Sec. 3, as well as those reviewed in Sec. 5.
It is however important to emphasize that the bulk here is fictitious and a spin-1/2 chain is perfectly well-defined without the bulk. One may view the bulk as merely a mathematical device to characterize the anomaly. On the other hand, it is also certainly possible to study the boundary of an actual crystalline SPT state. Although the physical contexts are different, the anomaly has the same implications for low-energy physics.
@@ 7 DISCUSSION
  • What is a fully general microscopic definition of anomalies? Currently, such definitions for anomalies of internal and translation symmetries have been discussed [37, 12, 38, 15, 39, 40, 41, 42], but the extension to include point-group symmetry is still largely unavailable. It is also important to generalize these definitions to continuum systems, such as Landau levels.
@@
  • ’t Hooft anomalies are originally defined as obstructions to gauging symmetries. What do gauge fields of spatial symmetries mean? Ref. [54] laid down the foundation of this subject at the level of certain coarse grained effective theories. Can such gauge fields be formulated more microscopically, based on which one can study the renormalization group flow of the properties of the spatial symmetry defects? These questions are also related to the point that the anomaly matching framework discussed in Sec. 4 relies on the crystalline equivalence principle, and it is desirable to develop a first-principle framework of anomaly matching.
@@ ACKNOWLEDGMENTS
LZ thanks Meng Guo, Yin-Chen He, Chong Wang, Jinmin Yi, Shiyu Zhou, and especially Ruizhi Liu and Weicheng Ye for collaborations on topics related to this review. MC thanks Maissam Barkeshli and Nathan Seiberg for collaborations on related topics. LZ is supported by the National University of Singapore start-up grants A-0009991-00-00 and A-0009991-01-00. MC is partially supported by NSF grant DMR-2424315.

### 29 Apr 2026 v2

@@ 4.4 Topological Quantum Spin Liquids
For instance, a $\mathbb{Z}_{2}$ TQSL has 4 types of anyons, often denoted by 1, $e$, $m$ and $\epsilon$, where 1 is a trivial anyon that is similar to the conventional magnons, $e$ and $m$ are both anyons with bosonic self-statistics but semionic mutual braiding statistics [^6] That two particles have semionic mutual braiding statistics means that when one particle moves around the other by a full circle, the many-body state will get a 1-1 phase factor., and $\epsilon$ can be viewed as the bound state of $e$ and $m$, which has fermionic self-statistics and semionic mutual braiding statistics with both $e$ and $m$ [78]. If the $\mathbb{Z}_{2}$ TQSL further has an $\mathrm{SO}(3)$ spin rotational symmetry, then nontrivial anyons $e$, $m$ and $\epsilon$ may carry half-integer-spin projective representations of the $\mathrm{SO}(3)$ symmetry, instead of only carrying integer-spin linear representations as do magnons. Anyons carrying half-integer spins is an example of symmetry fractionalization, and different patterns of symmetry fractionalization signal different symmetry-enriched $\mathbb{Z}_{2}$ TQSLs [79, 61].
@@
Below we discuss a concrete example, i.e., a $\mathbb{Z}_{2}$ TQSL in triangular-lattice spin-1/2 systems with $G_{\rm UV}=p6m\times\mathrm{SO}(3)\times\mathbb{Z}_{2}^{\mathsf{T}}$, first worked out in Refs. [84, 85]. In this context, we only need the fractionalization patterns of the $G_{\rm UV}$ symmetry on the $e$ and $m$ anyons, which can be largely understood using general arguments, without doing complicated calculations.
@@
We can further fix the fractionalization pattern of the $p6m$ symmetry on $m$ completely. To this end, imagine inserting an $\mathrm{SO}(3)$ monopole into the system, which can be viewed as a $2\pi$ flux of the $\mathrm{U}(1)$ spin rotation symmetry around the $z$ axis. As the $2\pi$ flux is gauge-equivalent to no flux, this process nucleates an anyon $v$. The fact that $e$ ($m$) carries half-integer (integer) spin means that $e$ ($m$) experiences a 1-1 (1) Aharonov-Bohm phase factor when circling around this flux, which uniquely fixes the braiding phase between $e$ ($m$) and $v$, and further fixes $v=m$. This argument is essentially Laughlin’s flux-threading argument in the fractional quantum Hall effect.
Finally, note that the discussion in Sec. 4.2 shows that the projective representation of the $p6m$ symmetry carried by the $\mathrm{SO}(3)$ monopole is completely fixed by the type of the lattice. Therefore, the projective representation under the $p6m$ symmetry carried by $m$ is also completely fixed by the type of the lattice, which can be read off from Sec. II B 1 in Ref. [31]. For example, for $m$, translations along the two directions do not commute.
Taken together, the above discussion shows that the symmetry fractionalization pattern of the entire $G_{\rm UV}=p6m\times\mathrm{SO}(3)\times\mathbb{Z}_{2}^{\mathsf{T}}$ symmetry on $m$ is fully fixed, and $e$ carries half-integer spin under $\mathrm{SO}(3)$ and Kramers doublet under $\mathbb{Z}_{2}^{\mathsf{T}}$. However, the fractionalization pattern of the $p6m$ symmetry on $e$ has not been determined. To determine it, one way to proceed is to perform the systematic anomaly-matching analysis in Ref. [90]. A general framework to calculate the pullback of the IR anomaly $\varphi^{*}(\omega_{\rm IR})$ of a symmetry-enriched TQSL is given in Ref. [91]. Using the UV anomaly obtained in Ref. [31], one can then classify all possible symmetry-enrichment patterns compatible with the anomaly-matching condition Eq. (18). We refer readers to these references for more details, as they are relatively technical. Alternatively, one can also derive the symmetry fractionalization pattern of $e$ using the physical picture that the background anyon ($e$ here) must exhibit the same fractionalization pattern of space group as the DOF in the unit cell (see Ref. [85]).
@@ 4.5 Compressible Phases
As our last example, we briefly review the matching of filling anomaly in compressible phases. Here we consider a system with $\mathrm{U}(1)$ and translation symmetries, and a compressible phase is one where the low-energy effective field theory of the system takes the same form at different filling fractions. The most familiar examples are superfluids and Fermi liquids (FLs). Being compressible is a rather strong condition. For example, Dirac and Weyl semimetals are not compressible in this sense, because when the filling fraction changes, they will become FLs and the effective theory changes.
By matching the filling anomaly, Ref. [92] pointed out that, in spatial dimensions larger than 1, the internal symmetry of the effective theory of a compressible phase cannot be a 0-form symmetry described by a finite-dimensional compact Lie group. To get some intuition of this result, note that the effective theory of a $(d+1)$-dimensional superfluid actually has a (d1)(d-1)-form symmetry, which corresponds to the conservation of vorticity [60, 93]. In FLs, the symmetry of the effective theory is a “loop group” of the $\mathrm{U}(1)$, which corresponds to the conservation of fermion number at each point on the Fermi surface [92]. In both cases, the symmetries of effective theories are not 0-form symmetries described by a finite-dimensional compact Lie group. We refer the readers to Refs. [92, 94] for more details. The above result leads to multiple further applications. For example, it can be used to show the Luttinger theorem [92], which was first obtained perturbatively [95] and later proved using non-perturbative topological arguments [96].
A more exotic type of compressible phase is the fractionalized Fermi liquids (FL) [97, 98], which can be viewed as an FL stacked on top of a TQSL discussed in Sec. $4.$ The anomaly of an FL can be viewed as the “sum” of those of an FL and a TQSL, and the anomaly matching condition imposes strong constraints on how the Luttinger theorem can be violated and which types of FL can emerge at a given filling factor [99, 100].
@@ 5 OTHER GENERALIZATIONS
In this section, we discuss various other generalizations of the LSM anomalies. There are also generalizations not covered in this review due to length constraints, and we refer the readers to Refs. [21, 22, 23, 24, 25, 26] for details.
@@ 5.1 Disordered LSM and Average Anomaly
In real solids, lattice symmetries are always broken by impurities. However, viewed as an ensemble (assuming self-averaging), lattice symmetries like translations are usually restored “on average”. That is, while the symmetry is completely broken for any given disorder realization, in the ensemble any two disorder realizations related to each other via translations occur with the same probability. Therefore, it is important to understand to what extent the LSM constraints apply in this case. Remarkably, LSM anomalies can still be robust if the lattice symmetries are only preserved on average [49, 11, 101, 102].
As a concrete example, consider a spin chain with an average translation symmetry, which still allows one to define a unit cell. Moreover, suppose there is an on-site $\mathrm{SO}(3)$ spin rotational symmetry such that each unit cell contains a spin-1/2 moment. If the translation symmetry was exact, in Sec. 2 we see that this symmetry setting would lead to a nontrivial LSM anomaly that forces all symmetric states to be LRE. Interestingly, when the translation symmetry is only preserved on average in an ensemble of such systems [^7] By the Imry-Ma argument, such an average translation symmetry cannot be broken spontaneously in 1D., then the probability for $\mathrm{SO}(3)$ symmetric ground states to be LRE approaches 1 in the thermodynamic limit [11]. That is, the ensemble is LRE. One can further establish an analog of Theorem 2.3 that bounds the decay of the ensemble-averaged spin correlation function [49].
@@ 5.2 Fermionic Systems
A physically clear example is spin-1/2 electrons at half filling (i.e., one electron per unit cell). Suppose there is a large Hubbard-like interaction suppressing double occupancy. Then the effective description of the system is a spin-1/2 system satisfying the previously discussed LSM theorem. It is then natural to expect that a similar constraint applies even when the single-occupancy constraint is not strictly enforced. That is, the system cannot be in a symmetric SRE state at half filling.
@@ 5.3 SPT-LSM Theorems
Other examples of SPT-LSM theorems are considered in Refs. [17, 18, 19]. For the general theory of SPT-LSM theorems, we refer the readers to Refs. [20, 8].
@@ 6 CRYSTALLINE SYMMETRY-PROTECTED TOPOLOGICAL PHASES
Finally, we discuss a fruitful approach to characterize ’t Hooft anomalies, the “anomaly inflow”. In this approach, the $d$-dimensional system with the anomalous symmetry is viewed as the boundary of a $(d+1)$-dimensional bulk, such that the whole system has no anomaly. That is, the non-gauge-invariant response of the boundary system to the background gauge field is cancelled by the response of the bulk. To achieve this, the bulk has to be a nontrivial SPT phase. This picture can be made precise in the QFT framework, establishing a one-to-one correspondence between ’t Hooft anomalies in $d$ dimensions and SPT phases in $(d+1)$ dimensions.
This anomaly inflow picture can also be applied to LSM anomalies, where the bulk is an SPT phase protected by internal and/or crystalline symmetries (hence “crystalline SPTs”) [27, 105]. As a concrete example, let us revisit the LSM anomaly for the spin-1/2 chain discussed in Sec. 2.1. The 2D bulk is easily seen to be a decoupled stack of Haldane chains (i.e., the 1D nontrivial SPT phase protected by $\mathrm{SO}(3)$ symmetry), each, say, along the $y$ direction. Moreover, these Haldane chains have a translation symmetry along the $x$ direction perpendicular to $y$. The spin-1/2 chain emerges at the edge along the $x$ direction. Similar constructions apply to other LSM anomalies discussed in Sec. 3, as well as those reviewed in Sec. 5. In all these examples, the anomalies are determined by the nature of the degrees of freedom and the symmetry actions on them, which must be compatible with the bulk properties. However, the filling anomalies in Sec. 2.4 and Sec. 4.5, which are due to the $\mathrm{U}(1)$ and translation symmetries, also depend on the eigenspace of the symmetry operator in which the system lies. For example, although the filling anomaly exists in subspaces with fractional filling factors, it disappears in subspaces with integer filling factors. In contrast, Theorem 2.4 applies to all symmetric states. For this reason, the filling anomalies do not admit a standard anomaly-inflow picture. In fact, there is no nontrivial SPT phase protected by both $\mathrm{U}(1)$ and translation symmetries in any dimension [92].
It is, however, important to emphasize that the bulk here is fictitious and a spin-1/2 chain is perfectly well-defined without the bulk. One may view the bulk as merely a mathematical device to characterize the anomaly. On the other hand, it is also certainly possible to study the boundary of an actual crystalline SPT state. Although the physical contexts are different, the anomaly has the same implications for low-energy physics.
@@ 7 DISCUSSION
  • What is a fully general microscopic definition of anomalies? Currently, such definitions for anomalies of internal and translation symmetries have been discussed [40, 12, 41, 15, 42, 43, 44, 45], but the extension to include point-group symmetry is still largely unavailable. It is also important to generalize these definitions to continuum systems, such as particles moving in a 2D continuum with a perpendicular magnetic field that leads to Landau levels.
@@
  • ’t Hooft anomalies are originally defined as obstructions to gauging symmetries. What do gauge fields of spatial symmetries mean? Ref. [59] laid down the foundation of this subject at the level of certain coarse-grained effective theories. Can such gauge fields be formulated more microscopically, based on which one can study the renormalization group flow of the properties of the spatial symmetry defects? These questions are also related to the point that the anomaly matching framework discussed in Sec. 4 relies on the crystalline equivalence principle, and it is desirable to develop a first-principles framework of anomaly matching.
@@ ACKNOWLEDGMENTS
LZ thanks Meng Guo, Yin-Chen He, Chong Wang, Jinmin Yi, Shiyu Zhou, and especially Ruizhi Liu and Weicheng Ye for collaborations on topics related to this review. MC thanks Maissam Barkeshli, Yang Qi and Nathan Seiberg for collaborations on related topics. LZ is supported by the National University of Singapore start-up grants A-0009991-00-00 and A-0009991-01-00. MC is supported by NSF grant DMR-2424315. This manuscript benefited from language improvements assisted by ChatGPT 5.